An ideal gas at 27 0C is compressed adiabatically to 8/27 of its original volume. if =5/3, then the rise in temperature is
 
a. 450 0C                                          b. 375 0C
 
c. 225 0C                                          d. 405 0C


Grade:9

3 Answers

Abhishek
7 years ago
375 C ixffg*+₹₹66₹(₹+_-"-₹7_&-₹7647)_+_7)4)_7425#6₹7)_()34217_867(₹4##;68₹6₹+"(₹-"++6_+&9-)&8587252&@-?₹+(-(`_74(`(₹5@38/_+₹??"-
Hamza
6 years ago
Since it is a adiabaric processT×V^(gamma-1)Gamma for a monoatomic gas=5/3T°V°^(2/3)=T`(8V°/27)^(2/3)So T`=675 Kand change in temp=(675-300)=375K
Kushagra Madhukar
3 years ago
Dear student,
Please find the solution to your problem.
 
Since it is a adiabaric process
Hence, TVγ-1 = constant
Where, γ = 5/3
Now,
T°V°2/3 = T’(8V°/27)2/3
So T’ = 675 K
Hence, change in temp = (675 – 300) = 375K
Hence, option b) is correct.
 
Thanks and regards,
Kushagra